OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through point (-3, 4) and perpendicular to a line that has slope 2/5

OpenStudy (anonymous):

(y-4)=2/5(x+3) (equation in point slope) then take the opposite recirpical of the slope: -5/2 and plug back into the equation (y-4)=(-5/2)(x+3) not sure what standard form is considered in your class but you can probably derive it from the equation above